Suppose we have a 1Gbps network connection from Boston to New York. There is a TCP link with window size is 6000B over the network. The one-way delay on the link it 25 ms.
Answer #1

Time taken to transmit 6000B over the network:

= (6000 * 8 bits) / (1 * 109 Gbps) = 48 * 10-6 sec = 0.048 ms

Hence, total delay = 25 + 0.048 = 25.048 ms